OpenStudy (anonymous):

Evaluate the expression below when x = 3. -2x - 8 -(-3x + 8)

OpenStudy (mathstudent55):

That means replace x with 3 and evaluate. Make sure to use the order of operations. -2x - 8 -(-3x + 8) -2(3) - 8 - [-3(3) + 8]

OpenStudy (anonymous):

i dont understand?

OpenStudy (mathstudent55):

-2x - 8 -(-3x + 8) -2(3) - 8 - [-3(3) + 8] = -6 - 8 - (-9 + 8) = -14 - (-1) = -14 + 1 = -13